Mayo is 37 years and Abdul is 6 years. how many years ago was the product of their ages 96?
Answer by Alan3354(69443)

Mayo is 37 years and Abdul is 6 years. how many years ago was the product of their ages 96?
==============
The difference is 31 years.
1*32 = 32
2*33 = 66
3*34 = 102
---------
There's no integer solution.